# Tighter Confidence Bounds for Sequential Kernel Regression

### v1

@@ Abstract
Confidence bounds are an essential tool for rigorously quantifying the uncertainty of predictions. In this capacity, they can inform the exploration-exploitation trade-off and form a core component in many sequential learning and decision-making algorithms. Tighter confidence bounds give rise to algorithms with better empirical performance and better performance guarantees. In this work, we use martingale tail bounds and finite-dimensional reformulations of infinite-dimensional convex programs to establish new confidence bounds for sequential kernel regression. We prove that our new confidence bounds are always tighter than existing ones in this setting. We apply our confidence bounds to the kernel bandit problem, where future actions depend on the previous history. When our confidence bounds replace existing ones, the KernelUCB (GP-UCB) algorithm has better empirical performance, a matching worst-case performance guarantee and comparable computational cost. Our new confidence bounds can be used as a generic tool to design improved algorithms for other kernelised learning and decision-making problems.
Kernels, bandits, confidence sequences, martingales, convex optimisation
@@ 1 Introduction
A confidence sequence for an unknown ground-truth function $f^{*}$ is a sequence of sets $\mathcal{F}_{1},\mathcal{F}_{2},\dots$, where each set $\mathcal{F}_{t}$ contains all functions $f$ that could plausibly be the ground-truth $f^{*}$, given all data available at time step $t$. For any point $x$, the corresponding upper confidence bound, $\max_{f\in\mathcal{F}_{t}}\{f(x)\}$, is thus the largest value that $f^{*}(x)$ could plausibly take, given all data available at time step $t$.
@@
The uncertainty estimates provided by confidence bounds are exceptionally useful in sequential learning and decision-making problems, such as bandits (Lattimore & Szepesvári, 2020) and reinforcement learning (Sutton & Barto, 2018), where they can be used to design exploration strategies (Auer, 2002), establish stopping criteria (Jamieson et al., 2014) and guarantee safety (Sui et al., 2015; Berkenkamp et al., 2017). In general, tighter confidence bounds give rise to sequential learning and decision-making algorithms with better empirical performance and better performance guarantees. As many modern algorithms use uncertainty estimates for flexible nonparametric function approximators, such as kernel methods, it is vital to establish tight confidence bounds for infinite-dimensional function classes.
In this work, we develop a new confidence sequence for sequential kernel regression. The corresponding upper confidence bound at any point $x$ is the solution of an infinite-dimensional convex program in the kernel Hilbert space. We show that this can be reformulated using either a representer theorem or Lagrangian duality, yielding finite-dimensional convex programs, which are efficiently solvable and certifiable. We propose several methods for computing our confidence bounds, which can replace the confidence bounds used in kernel bandit algorithms, such as KernelUCB (Valko et al., 2013). We demonstrate that the resulting kernel bandit algorithms have better empirical performance, matching worst-case performance guarantees and comparable computational cost. Finally, we prove that our new confidence bounds are always tighter than existing confidence bounds in the same setting (see Fig. 1).
@@
Figure 1: Tighter confidence bounds for sequential kernel regression. The upper and lower confidence bounds used by our Kernel CMM-UCB method (left), the confidence bounds from (Abbasi-Yadkori, 2012) (AY-GP-UCB) (middle-right), and the confidence bounds used by Improved GP-UCB (IGP-UCB) (Chowdhury & Gopalan, 2017) (right) for a Matérn kernel test function with smoothness $\nu=3/2$ and lengthscale $\ell=0.5$. Our Kernel CMM-UCB method – and its relaxed versions DMM-UCB and AMM-UCB – produces confidence bounds that are visibly closer to the ground-truth function (dashed line) than those of AY-GP-UCB and IGP-UCB (cf. also Sec. 6.1).
@@ 2 Problem Statement
$y_{t}=f^{*}(x_{t})+\epsilon_{t}.$
$f^{*}\in\mathcal{H}$ is an unknown function in a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ and $\epsilon_{1},\epsilon_{2},\dots$ are noise variables. We allow the covariates to be generated in an arbitrary sequential fashion, meaning each $x_{t}$ can depend on $x_{1},y_{1},\dots,x_{t-1},y_{t-1}$. We assume that the RKHS norm of $f^{*}$ is bounded by $B$, i.e. $\left\|f^{*}\right\|_{\mathcal{H}}\leq B$, and that, conditioned on $x_{1},y_{1},\dots,x_{t-1},y_{t-1},x_{t}$, each $\epsilon_{t}$ is $\sigma$-sub-Gaussian.
Our first aim is to construct a confidence sequence for the function $f^{*}$, which we now define formally. For any level $\delta\in(0,1]$, a $(1-\delta)$-confidence sequence is a sequence $\mathcal{F}_{1},\mathcal{F}_{2},\dots$ of subsets of $\mathcal{H}$, such that each $\mathcal{F}_{t}$ can be calculated using the data $x_{1},y_{1},\dots,x_{t},y_{t}$ and the sequence satisfies the coverage condition
@@
In words, this coverage condition says that with high probability, $f^{*}$ lies in $\mathcal{F}_{t}$ for all times $t\geq 1$ simultaneously. We call each set in a confidence sequence a confidence set. Our second aim is to construct confidence bounds for the value of $f^{*}(x)$ at every $x\in\mathcal{X}$. For any level $\delta\in(0,1]$, an anytime-valid $(1-\delta)$-upper confidence bound (UCB) is a sequence $u_{1},u_{2},\dots$ of functions on $\mathcal{X}$, such that each $u_{t}$ can be calculated using the data $x_{1},y_{1},\dots,x_{t},y_{t}$ and the sequence satisfies
@@
This condition says that with high probability, $u_{t}(x)$ is an upper bound on $f^{*}(x)$ for all $t\geq 1$ and all $x\in\mathcal{X}$ simultaneously. The definition of an anytime-valid lower confidence bound (LCB) is analogous. If we already have a confidence sequence, we can use it to construct upper and lower confidence bounds. In particular, if $\mathcal{F}_{1},\mathcal{F}_{2},\dots$ is a $(1-\delta)$-confidence sequence, then with probability at least $1-\delta$, for all $t\geq 1$ and all $x\in\mathcal{X}$, we have
$\min_{f\in\mathcal{F}_{t}}\left\{f(x)\right\}\,\leq\,f^{*}(x)\,\leq\,\max_{f\in\mathcal{F}_{t}}\left\{f(x)\right\}.$ (1)
@@
For any positive integer $n$, $[n]:=\{1,\dots,n\}$. We use matrix notation to denote the inner product $f_{1}^{\top}f_{2}:=\left\langle f_{1},f_{2}\right\rangle_{\mathcal{H}}$ and outer product $f_{1}f_{2}^{\top}:=f_{1}\left\langle f_{2},\cdot\right\rangle_{\mathcal{H}}$ between two RKHS functions $f_{1},f_{2}\in\mathcal{H}$. We denote the kernel function by $k(\cdot,\cdot)$. For a sequence of observations $x_{1},\dots,x_{t}$, we define $\boldsymbol{k}_{t}(x):=[k(x,x_{1}),\dots,k(x,x_{t})]^{\top}$ and $\Phi_{t}:=[k(x_{1},\cdot),\dots,k(x_{t},\cdot)]^{\top}$. We let $\boldsymbol{K}_{t}:=\{k(x_{i},x_{j})\}_{i\in[t],j\in[t]}=\Phi_{t}\Phi_{t}^{\top}$ denote the kernel matrix and $\boldsymbol{V}_{t}:=\sum_{s=1}^{t}k(\cdot,x_{s})k(\cdot,x_{s})^{\top}=\Phi_{t}^{\top}\Phi_{t}$ the empirical covariance operator. For a parameter $\alpha\geq 0$, we define
@@ 3 Related Work
Several confidence sequences/bounds have been proposed for the sequential kernel regression problem that we consider. The confidence bounds in (Srinivas et al., 2010) are perhaps the most well-known. The two most widely used, which we compare ours against, were developed by Abbasi-Yadkori et al. (2011); Abbasi-Yadkori (2012) and Chowdhury & Gopalan (2017), using concentration inequalities for self-normalised processes (de la Peña et al., 2004, 2009). Subsequently, Durand et al. (2018) derived empirical Bernstein-type versions of these confidence bounds, and Whitehouse et al. (2023) showed that for certain kernels, these confidence bounds can be tightened by regularising in proportion to the smoothness of the kernel. Confidence sequences for this setting have also been derived using online-to-confidence-set conversions (Abbasi-Yadkori et al., 2012; Abbasi-Yadkori, 2012).
With applications to Bayesian optimisation and kernel bandits in mind, Neiswanger & Ramdas (2021) and Emmenegger et al. (2023) developed confidence sequences using martingale tail bounds for sequential likelihood ratios (assuming Gaussian rather than sub-Gaussian noise). However, cumulative regret bounds for KernelUCB-style bandit algorithms using these confidence bounds are not established in (Neiswanger & Ramdas, 2021), and the regret bound in (Emmenegger et al., 2023) is worse than the one obtained with the confidence bounds in (Abbasi-Yadkori, 2012) or with our new confidence bounds.
The work most closely related to ours is (Flynn et al., 2023), which proposes new confidence bounds for linear bandits, using martingale tail bounds and the method of mixtures. While our Eq. (8) is a simple kernel generalisation of their Eq. (5), deriving confidence bounds is more involved in our kernel setting, because $\max_{f\in\mathcal{F}_{t}}\{f(x)\}$ is now an infinite-dimensional optimisation problem. In their linear (i.e. finite-dimensional) setting, Flynn et al. (2023) bound the regret of their algorithms in terms of the feature vector dimension. The feature dimension is generally infinite in the kernel setting, so we instead bound the regret in terms of the maximum information gain (Srinivas et al., 2010).
@@ 4 Confidence Bounds for Kernel Regression
In this section, we use a tail bound for martingale mixtures from (Flynn et al., 2023) to develop tighter confidence sequences and bounds for sequential kernel regression.
@@ 4.1 Martingale Mixture Tail Bounds
[…]
@@
Flynn et al. (2023) show that $M_{t}(\boldsymbol{g}_{t},\boldsymbol{\lambda}_{t})$ is a non-negative martingale, and that, due to conditions (a)-(c) on the mixture distributions, the mixture $\mathbb{E}_{\boldsymbol{g}_{t}\sim P_{t}}\left[M_{t}(\boldsymbol{g}_{t},\boldsymbol{\lambda}_{t})\right]$ is also a non-negative martingale. This means that Ville’s inequality (Ville, 1939) can be applied. Thm. 5.1 in (Flynn et al., 2023) provides a time-uniform tail bound for the martingale mixture $\mathbb{E}_{\boldsymbol{g}_{t}\sim P_{t}}\left[M_{t}(\boldsymbol{g}_{t},\boldsymbol{\lambda}_{t})\right]$.
Theorem 4.1 (Tail Bound for Adaptive Martingale Mixtures (Theorem 5.1 of (Flynn et al., 2023)).
For any $\delta\in(0,1)$, any adaptive sequence of mixture distributions $(P_{t}|t\in\mathbb{N})$, and any sequence of predictable random variables $(\lambda_{t}|t\in\mathbb{N})$, it holds with probability at least $1-\delta$ that
@@
We choose $(\mathcal{D}_{t}|t\in\mathbb{N})$ to be any filtration such that $y_{t}$ and $x_{t+1}$ are both $\mathcal{D}_{t}$-measurable, e.g. $\mathcal{D}_{t}=\sigma(x_{1},y_{1},\dots,x_{t},y_{t},x_{t+1})$. Focusing on linear bandits, Flynn et al. (2023) choose $Z_{t}(g_{t})=(g_{t}-\phi(x_{t})^{\top}\boldsymbol{\theta}^{*})\epsilon_{t}$, where $\phi(\cdot)^{\top}\boldsymbol{\theta}^{*}$ is a linear reward function, and show that under this choice, Eq. (6) reduces to a convex constraint for the finite-dimensional parameter vector $\boldsymbol{\theta}^{*}$. In our kernel setting, we choose $Z_{t}(g_{t})=(g_{t}-f^{*}(x_{t}))\epsilon_{t}$, leading to a convex constraint for the function $f^{*}$. As $Z_{t}(g_{t})$ is linear in the noise variable $\epsilon_{t}$, $\psi_{t}(g_{t},\lambda_{t})$ can be upper bounded using the $\sigma$-sub-Gaussian property of $\epsilon_{t}$.
$\psi_{t}(g_{t},\lambda_{t})\,\leq\,\tfrac{1}{2}\sigma^{2}\lambda_{t}^{2}(g_{t}-f^{*}(x_{t}))^{2}.$
Using this upper bound, choosing $\lambda_{t}\equiv 1/\sigma^{2}$, and rearranging Eq. (6) (see App. A.1), we obtain an upper bound on the Euclidean distance between the vector of ground-truth function values $\boldsymbol{f}_{t}^{*}:=[f^{*}(x_{1}),\dots,f^{*}(x_{t})]^{\top}$ and the observed response vector $\boldsymbol{y}_{t}:=[y_{1},\dots,y_{t}]^{\top}$.
$\begin{split}&\|\boldsymbol{f}_{t}^{*}-\boldsymbol{y}_{t}\|_{2}^{2}~{}\leq~{}2\sigma^{2}\ln(1/\delta)\\ &\quad-2\sigma^{2}\ln\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\|\boldsymbol{g}_{t}-\boldsymbol{y}_{t}\|_{2}^{2}\right)\right]\right).\end{split}$ (7)
Next, we choose the mixture distributions $P_{t}=\mathcal{N}(\boldsymbol{0},c\boldsymbol{K}_{t})$, resembling a zero-mean Gaussian process with covariance scaled by $c>0$. There are several reasons for this choice of $P_{t}$. First, for an appropriate covariance scale $c>0$, we prove that the confidence bounds obtained from (7) with $P_{t}=\mathcal{N}(\boldsymbol{0},c\boldsymbol{K}_{t})$ are always tighter than those in (Abbasi-Yadkori, 2012) and (Chowdhury & Gopalan, 2017) (see Sec. 6.1). Second, any Gaussian $P_{t}$ yields a convenient closed-form expression for the expected value in (7) (see App. A.2). Substituting this expression into (7), we obtain
$\|\boldsymbol{f}_{t}^{*}-\boldsymbol{y}_{t}\|_{2}^{2}~{}\leq$ $~{}\,\boldsymbol{y}_{t}^{\top}\left(\boldsymbol{I}+\frac{c\boldsymbol{K}_{t}}{\sigma^{2}}\right)^{-1}\boldsymbol{y}_{t}$ (8)
$+\sigma^{2}\ln\det\left(\boldsymbol{I}+\frac{c\boldsymbol{K}_{t}}{\sigma^{2}}\right)+2\sigma^{2}\ln\frac{1}{\delta}$
@@
This inequality states that with probability at least $1-\delta$, at every step $t$, the vector of ground-truth function values $\boldsymbol{f}_{t}^{*}$ lies within a sphere of radius $R_{t}$ around the observed response vector $\boldsymbol{y}_{t}$.

### v2

@@ Abstract
Confidence bounds are an essential tool for rigorously quantifying the uncertainty of predictions. They are a core component in many sequential learning and decision-making algorithms, with tighter confidence bounds giving rise to algorithms with better empirical performance and better performance guarantees. In this work, we use martingale tail inequalities to establish new confidence bounds for sequential kernel regression. Our confidence bounds can be computed by solving a conic program, although this bare version quickly becomes impractical, because the number of variables grows with the sample size. However, we show that the dual of this conic program allows us to efficiently compute tight confidence bounds. We prove that our new confidence bounds are always tighter than existing ones in this setting. We apply our confidence bounds to kernel bandit problems, and we find that when our confidence bounds replace existing ones, the KernelUCB (GP-UCB) algorithm has better empirical performance, a matching worst-case performance guarantee and comparable computational cost.
@@ 1 INTRODUCTION
A confidence sequence for an unknown function $f^{\star}$ is a sequence of sets $\mathcal{F}_{1},\mathcal{F}_{2},\dots$, where each set $\mathcal{F}_{t}$ contains all functions that could plausibly be $f^{\star}$, given all data available at time step $t$. For any point $x$, the corresponding upper confidence bound, $\max_{f\in\mathcal{F}_{t}}\{f(x)\}$, is thus the largest value that $f^{\star}(x)$ could plausibly take, given all data available at time step $t$.
@@
The uncertainty estimates provided by confidence bounds are exceptionally useful in sequential learning and decision-making problems, such as bandits (Lattimore and Szepesvári,, 2020) and reinforcement learning (Sutton and Barto,, 2018), where they can be used to design exploration strategies (Auer,, 2002), establish stopping criteria (Jamieson et al.,, 2014) and guarantee safety (Sui et al.,, 2015; Berkenkamp et al.,, 2017). In general, tighter confidence bounds give rise to sequential learning and decision-making algorithms with better empirical performance and better performance guarantees. As many modern algorithms use uncertainty estimates for flexible nonparametric function approximators, such as kernel methods, it is vital to establish tight confidence bounds for infinite-dimensional function classes.
In this work, we develop a new confidence sequence for sequential kernel regression. The corresponding upper confidence bound at any point $x$ is the solution of an infinite-dimensional convex program in the kernel Hilbert space. This can be reformulated using the representer theorem, yielding a finite-dimensional convex program, and our Kernel CMM-UCB method. However, computing the solution becomes expensive, because the number of variables grows with the sample size. One solution is to use an analytic upper bound on the exact confidence bound, as we do in our Kernel AMM-UCB method. Unfortunately, this analytic confidence bound can be noticeably looser. A better solution is to compute confidence bounds by solving the Lagrangian dual of the original convex program, which can be done cheaply. This results in our Kernel DMM-UCB method, which achieves the best of both worlds: cheap and tight confidence bounds (see Fig. 1).
@@
Figure 1: The upper and lower confidence bounds used by our Kernel CMM-UCB method (left), the confidence bounds from (Abbasi-Yadkori,, 2012) (AY-GP-UCB) (middle-right), and the confidence bounds used by Improved GP-UCB (IGP-UCB) (Chowdhury and Gopalan,, 2017) (right) for a Matérn kernel test function with smoothness $\nu=3/2$ and lengthscale $\ell=0.5$. Our Kernel CMM-UCB method – and its relaxed versions DMM-UCB and AMM-UCB – produces confidence bounds that are visibly closer to the ground-truth function (dashed line) than those of AY-GP-UCB and IGP-UCB (cf. also Sec. 6.1).
Our new confidence bounds can be used in kernel bandit algorithms, such as KernelUCB/GP-UCB. We show that using our confidence bounds preserves the worst-case performance guarantees of KernelUCB. In doing so, we refine the standard KernelUCB regret analysis by using new bounds on the number of large elliptical potentials. With these new tools, we show that in the noiseless setting, KernelUCB with our confidence bounds enjoys a $\mathrm{polylog}(T)$ cumulative regret bound when the RBF kernel is used, and a sub-$\sqrt{T}$ regret bound when the Matérn kernel with smoothness $\nu\geq d/\sqrt{2}$ is used. Finally, we demonstrate that the empirical performance of KernelUCB can be vastly improved by using our new confidence bounds.
@@ 2 PROBLEM STATEMENT
$y_{t}=f^{\star}(x_{t})+\epsilon_{t}.$
$f^{\star}\in\mathcal{H}$ is an unknown function in a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ and $\epsilon_{1},\epsilon_{2},\dots$ are noise variables. We allow the covariates to be generated in an arbitrary sequential fashion, meaning each $x_{t}$ can depend on $x_{1},y_{1},\dots,x_{t-1},y_{t-1}$. We assume that the RKHS norm of $f^{\star}$ is bounded by $B$, i.e. $\left\|f^{\star}\right\|_{\mathcal{H}}\leq B$. For most of the paper, we assume that, conditioned on $x_{1},y_{1},\dots,x_{t-1},y_{t-1},x_{t}$, each $\epsilon_{t}$ is a $\sigma$-sub-Gaussian random variable. In Sec. 6.3, we consider the noiseless setting, where $\epsilon_{t}\equiv 0$, and we observe $y_{t}=f^{\star}(x_{t})$.
Our first aim is to construct a confidence sequence for the function $f^{\star}$, which we now define formally. For any level $\delta\in(0,1]$, a $(1-\delta)$-confidence sequence is a sequence $\mathcal{F}_{1},\mathcal{F}_{2},\dots$ of subsets of $\mathcal{H}$, such that each $\mathcal{F}_{t}$ can be calculated using the data $x_{1},y_{1},\dots,x_{t},y_{t}$ and the sequence satisfies the coverage condition
@@
In words, this coverage condition says that with high probability (over the random draw of $x_{1},y_{1},x_{2},y_{2},\dots$), $f^{\star}$ lies in $\mathcal{F}_{t}$ for all times $t\geq 1$ simultaneously. We call each set in a confidence sequence a confidence set. Our second aim is to construct confidence bounds for the value of $f^{\star}(x)$ at every $x\in\mathcal{X}$. For any level $\delta\in(0,1]$, an anytime-valid $(1-\delta)$-upper confidence bound (UCB) is a sequence $u_{1},u_{2},\dots$ of functions on $\mathcal{X}$, such that each $u_{t}$ can be calculated using the data $x_{1},y_{1},\dots,x_{t},y_{t}$, and the sequence satisfies
@@
This condition says that with high probability, $u_{t}(x)$ is an upper bound on $f^{\star}(x)$ for all $t\geq 1$ and all $x\in\mathcal{X}$ simultaneously. The definition of an anytime-valid lower confidence bound (LCB) is analogous. If we already have a confidence sequence, we can use it to construct upper and lower confidence bounds. In particular, if $\mathcal{F}_{1},\mathcal{F}_{2},\dots$ is a $(1-\delta)$-confidence sequence, then with probability at least $1-\delta$, for all $t\geq 1$ and all $x\in\mathcal{X}$, we have
$\min_{f\in\mathcal{F}_{t}}\left\{f(x)\right\}\,\leq\,f^{\star}(x)\,\leq\,\max_{f\in\mathcal{F}_{t}}\left\{f(x)\right\}.$ (1)
@@
For any $n\in\mathbb{N}$, $[n]:=\{1,\dots,n\}$. We use matrix notation to denote the inner product $f_{1}^{\top}f_{2}:=\left\langle f_{1},f_{2}\right\rangle_{\mathcal{H}}$ and outer product $f_{1}f_{2}^{\top}:=f_{1}\left\langle f_{2},\cdot\right\rangle_{\mathcal{H}}$ between two functions $f_{1},f_{2}\in\mathcal{H}$. We denote the kernel function by $k(\cdot,\cdot)$. For a sequence of observations $x_{1},\dots,x_{t}$, we define $\boldsymbol{k}_{t}(x):=[k(x,x_{1}),\dots,k(x,x_{t})]^{\top}$ and $\Phi_{t}:=[k(x_{1},\cdot),\dots,k(x_{t},\cdot)]^{\top}$. We let $\boldsymbol{K}_{t}:=\{k(x_{i},x_{j})\}_{i\in[t],j\in[t]}=\Phi_{t}\Phi_{t}^{\top}$ denote the kernel matrix and $\boldsymbol{V}_{t}:=\sum_{s=1}^{t}k(\cdot,x_{s})k(\cdot,x_{s})^{\top}=\Phi_{t}^{\top}\Phi_{t}$ the empirical covariance operator. For $\alpha\geq 0$, we define
@@ 3 RELATED WORK
Several confidence sequences/bounds have been proposed for the sequential kernel regression problem that we consider. The confidence bounds in (Srinivas et al.,, 2010) are perhaps the most well-known. The two most widely used were developed by Abbasi-Yadkori et al., (2011); Abbasi-Yadkori, (2012) and Chowdhury and Gopalan, (2017), using concentration inequalities for self-normalised processes (de la Peña et al.,, 2004, 2009). Subsequently, Durand et al., (2018) derived empirical Bernstein-type versions of these confidence bounds, and Whitehouse et al., (2023) showed that for certain kernels, these confidence bounds can be tightened by regularising in proportion to the smoothness of the kernel. Confidence sequences for this setting have also been derived using online-to-confidence-set conversions (Abbasi-Yadkori et al.,, 2012; Abbasi-Yadkori,, 2012).
With applications Bayesian optimisation and kernel bandits in mind, Neiswanger and Ramdas, (2021) and Emmenegger et al., (2023) developed confidence sequences using martingale tail bounds for sequential likelihood ratios. However, cumulative regret bounds are not established in (Neiswanger and Ramdas,, 2021), and the regret bound in (Emmenegger et al.,, 2023) is worse than the regret bounds in (Abbasi-Yadkori,, 2012) and in this paper.
In the setting where $\epsilon_{1},\epsilon_{2},\dots$ are bounded deterministic errors, Scharnhorst et al., (2022) derived confidence bounds using convex programming and Lagrangian duality. However, these confidence bounds are not comparable to ours. In our setting, the random errors $\epsilon_{1},\epsilon_{2},\dots$ average out to 0, and our confidence bounds contract around $f^{\star}$. This is not the case for the confidence bounds in (Scharnhorst et al.,, 2022).
The work most closely related to ours is (Flynn et al.,, 2023), in which an improved confidence sequence for linear bandits is developed. While our Eq. (8) is a simple kernel generalisation of their Eq. (5), deriving confidence bounds is more involved in our kernel setting. In particular, $\max_{f\in\mathcal{F}_{t}}\{f(x)\}$ is now an infinite-dimensional optimisation problem, which makes our DMM-UCB method vital for balancing tightness and efficiency. Moreover, even in the simpler linear setting, the regret analysis in (Flynn et al.,, 2023) fails to provide a sub-$\sqrt{T}$ rate as $\sigma$ goes to zero.
@@ 4 CONFIDENCE BOUNDS FOR KERNEL REGRESSION
In this section, we describe our new confidence sequences and bounds for sequential kernel regression.
@@ 4.1 Martingale Mixture Tail Bounds
We use a martingale tail bound from (Flynn et al.,, 2023) to derive a data-dependent constraint for $f^{\star}$. We begin by describing the general setting in which it holds. There is a stochastic process $(Z_{t}(g)|t\in\mathbb{N},g\in\mathbb{R})$, indexed by a time $t$ and a real number $g$, which is adapted to a filtration $(\mathcal{D}_{t}|t\in\mathbb{N})$. In this paper, $(\mathcal{D}_{t}|t\in\mathbb{N})$ is any filtration such that $y_{t}$ and $x_{t+1}$ are both $\mathcal{D}_{t}$-measurable, e.g. $\mathcal{D}_{t}=\sigma(x_{1},y_{1},\dots,x_{t},y_{t},x_{t+1})$. In addition, there is a sequence of predictable “guesses” $(g_{t}\in\mathbb{R}|t\in\mathbb{N})$, and a sequence of predictable random variables $(\lambda_{t}\in\mathbb{R}|t\in\mathbb{N})$. We use the shorthand $\boldsymbol{g}_{t}=[g_{1},\dots,g_{t}]^{\top}$ and $\boldsymbol{\lambda}_{t}=[\lambda_{1},\dots,\lambda_{t}]^{\top}$. We define
@@
[…]
Theorem 4.1 (Theorem 5.1 of (Flynn et al.,, 2023)).
For any $\delta\in(0,1)$, any sequence of distributions $(P_{t}|t\in\mathbb{N})$ satisfying (a) and (b), and any sequence of predictable random variables $(\lambda_{t}|t\in\mathbb{N})$, it holds with probability at least $1-\delta$ that
@@
To turn the inequality in Thm. 4.1 into a constraint for $f^{\star}$, all that remains is to specify the process $Z_{t}(g)$, the random variables $\lambda_{t}$ and the distributions $P_{t}$. Intuitively, each $g_{t}$ is a guess for the value $y_{t}$, given the previous data $x_{1},y_{1},\dots,x_{t-1},y_{t-1},x_{t}$. We should therefore choose each $P_{t}$ such that (a) and (b) are satisfied and $P_{t}$ assigns high probability to the vector $[y_{1},\dots,y_{t}]^{\top}$. The choices of $Z_{t}(g)$ and $\lambda_{t}$ determine the shape of the resulting confidence sequence. We choose $Z_{t}(g)=(g-f^{\star}(x_{t}))\epsilon_{t}$, which leads to a convex quadratic constraint for $f^{\star}$. As $Z_{t}(g)$ is linear in the noise variable $\epsilon_{t}$, $\psi_{t}(g,\lambda_{t})$ can be upper bounded using the $\sigma$-sub-Gaussian property of $\epsilon_{t}$ (cf. Eq. (15)).
$\psi_{t}(g_{t},\lambda_{t})\,\leq\,\tfrac{1}{2}\sigma^{2}\lambda_{t}^{2}(g_{t}-f^{\star}(x_{t}))^{2}.$
Using this upper bound, choosing $\lambda_{t}\equiv 1/\sigma^{2}$, and rearranging Eq. (6) (see App. A.1), we obtain an upper bound on the squared Euclidean distance between the vector of function values $\boldsymbol{f}_{t}^{\star}:=[f^{\star}(x_{1}),\dots,f^{\star}(x_{t})]^{\top}$ and the response vector $\boldsymbol{y}_{t}:=[y_{1},\dots,y_{t}]^{\top}$.
$\begin{split}&\|\boldsymbol{f}_{t}^{*}-\boldsymbol{y}_{t}\|_{2}^{2}~{}\leq~{}2\sigma^{2}\ln(1/\delta)\\ &\quad-2\sigma^{2}\ln\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(-\tfrac{1}{2\sigma^{2}}\|\boldsymbol{g}_{t}-\boldsymbol{y}_{t}\|_{2}^{2}\right)\right]\right).\end{split}$ (7)
For the mixture distributions, we choose $P_{t}=\mathcal{N}(\boldsymbol{0},c\boldsymbol{K}_{t})$, resembling a zero-mean Gaussian process with covariance scaled by $c>0$. There are several reasons for this choice of $P_{t}$. First, for an appropriate covariance scale $c>0$, the confidence bounds obtained from (7) with $P_{t}=\mathcal{N}(\boldsymbol{0},c\boldsymbol{K}_{t})$ are always tighter than those in (Abbasi-Yadkori,, 2012) and (Chowdhury and Gopalan,, 2017) (see Sec. 6.1). Second, any Gaussian $P_{t}$ yields a convenient closed-form expression for the expected value in (7) (see App. A.2). Substituting this expression into (7), we obtain
$\|\boldsymbol{f}_{t}^{\star}-\boldsymbol{y}_{t}\|_{2}^{2}~{}\leq$ $~{}\,\boldsymbol{y}_{t}^{\top}\left(\boldsymbol{I}+\tfrac{c}{\sigma^{2}}\boldsymbol{K}_{t}\right)^{-1}\boldsymbol{y}_{t}$ (8)
$+\sigma^{2}\ln\det\left(\boldsymbol{I}+\tfrac{c}{\sigma^{2}}\boldsymbol{K}_{t}\right)+2\sigma^{2}\ln\frac{1}{\delta}$
@@
This inequality states that with probability at least $1-\delta$, at every step $t$, the vector of ground-truth function values $\boldsymbol{f}_{t}^{\star}$ lies within a sphere of radius $R_{t}$ around the observed response vector $\boldsymbol{y}_{t}$.